Derive an expression in the form p/q, where p and q are integers, for the recurring decimal 4.312121212......

Answer by rothauserc(4718)

let x = 4.312121212, then
100x = 431.121212
1000x = 43112.121212
1000x - 100x = 43112.121212 - 431.121212
9900x = 42681
x = 42681 / 9900